Differentiate the following function: f(x)=e^x+x^e
1
Expert's answer
2020-12-02T20:09:18-0500

We need to find derivative of function: "f(x)=e^x+x^e"

There is a sum of two elementary functions - power of x and simple exponential function.

If "e" is Euler's number, the derivative of "e^x" is "e^x". This is a property of the Euler number

And another function is exponentiation, so derivative is

"\frac{d(x^e)}{dx} = ex^{e-1}"

Using the summation property, we get answer

"\frac{d}{dx}(f(x))=\frac{d}{dx}(e^x+x^e)=\frac{d}{dx}(e^x)+\frac{d}{dx}(x^e)=e^x+ex^{e-1}"